# AS3D: 2D-Assisted Cross-Modal Understanding with Semantic-Spatial Scene Graphs for 3D Visual Grounding

### v2

@@ III-B Semantic-Spatial Scene Graph
In the 3D visual grounding task, the precise differentiation between identical type objects fundamentally hinges on positional context and referential relationships, as detailed in the given description. To tackle this issue, we introduce graph-based semantic-spatial modeling for relation learning within intricate scenes. Our framework employs the rotated bounding box encoding paradigm [9] to facilitate multi-view spatial comprehension via decoupled shared point cloud features. We augment semantic understanding by formulating a scene graph, which creates explicit correspondences between textual object types and multi-modal feature embeddings. To enhance spatial reasoning capacities, we integrate object-level spatial priors with global spatial encoding [13], incorporating them into the object features prior to graph attention operations. This architecture thereby enables comprehensive scene modeling of intra-object semantic relations and inter-object spatial dependencies.
III-B1 Graph Generation via Semantic Relations
@@
In complex scenes, the target description often entails multiple referential relationships. The innate capability of scene graphs to capture dependencies among various nodes aids in effectively processing intricate scene compositions. Prior studies [17, 16, 47] either construct fully-connected scene graphs or establish connections with neighboring nodes, implicitly learning the node and text embedding relationships. However, many reference orientations are not dependent on proximity, and a high number of objects can introduce uncertainties in node aggregation. We propose an explicit construction of graph node relationships using textual semantics that correspond to object features with multi-modal attributes. First, we match 3D feature representations with alignment training to class-aware prompts as described in Section III-A1. Then, we utilize the SpaCy lemmatization module to standardize the description text by converting words to their base forms, thereby significantly improving terminological consistency across different morphological variants. Finally, we use lexemes from the same semantic category in the modified description text to identify related objects. These matched objects, considered potential targets or references in localization cues, are then linked within the graph.
III-B2 Relation Learning via Graph Attention
@@
The graph nodes are input into a graph attention layer to facilitate multi-relational interactions, where feature representations are enhanced by aggregating heterogeneous relationship information from neighboring nodes. The $i$-th node value is updated through a multi-head attention mechanism:
@@
where $\mathcal{N}(i)$ denotes the set of neighboring nodes of vertex $i$, $A_{ij}\in\{0,1\}$ denotes the original adjacency relationship between nodes $i$ and $j$, $\alpha_{ij}^{k}\in(0,1]$ captures the normalized attention weight from node $j$ to $i$ through the $k$-th attention head. $W^{k}$ is the learnable projection matrix for the $k$-th head, $\sigma(\cdot)$ specifies the non-linear activation function.
TABLE I: Comparison with other recent methods on Nr3D and Sr3D
@@
AS3D (ours) - 67.2% 61.4% 73.3% 65.7% 68.0% 76.0% 71.7% 77.8% 67.2% 76.4%
vs. MiKASA [13] $\uparrow$2.8$\uparrow$2.0$\uparrow$3.6$\uparrow$0.3$\uparrow$4.0$\uparrow$0.8$\uparrow$4.4$\downarrow$0.8$\downarrow$3.2$\uparrow$1.0%
vs. MA2TransVG [12] $\uparrow$2.0$\uparrow$3.8$\uparrow$2.1$\uparrow$3.2$\uparrow$2.6$\uparrow$2.1$\uparrow$2.4$\uparrow$1.8$\uparrow$2.7$\uparrow$2.6%
TABLE II: Comparison with other recent methods on Scanrefer.
  Methods 2D Views Unique Multiple Overall
Num Enc
ScanRefer [20] - - 46.19% 21.26% 26.10%
TGNN [16] - - 56.80% 23.18% 29.70%
InstanceRefer [43] - - 66.83% 24.77% 32.93%
MVT [9] - - 66.45% 25.26% 33.26%
ViewRefer [10] - - - 26.50% 33.66%
3DVG-Trans [25] - - 58.47% 28.70% 34.47%
BUTD-DETR [50] - - 61.24% 32.81% 37.05%
3D-LP [52] - - 62.60% 32.18% 38.08%
EDA [49] - - 68.57% 37.64% 42.26%
MCLN [51] - - 68.36% 38.41% 42.64%
ScanRefer [20] all ENet 53.51% 21.11% 27.40%
3DVG-Trans [25] all ENet 60.64% 28.42% 34.67%
3D-SPS [27] all ENet 66.72% 29.82% 36.98%
3DJCG [29] all ENet 64.34% 30.82% 37.33%
D3Net [26] all ENet 72.04% 30.05% 37.87%
3D-LP [52] all ENet 64.61% 33.41% 39.46%
3DVLP [46] all ENet 70.04% 33.40% 40.51%
MA2TransVG [12] all ENet 74.1% 41.4% 45.7%
M3DRef-CLIP [48] 3 CLIP 77.2% 36.8% 44.7%
AS3D (ours) 1 CLIP 64.08% 40.87% 45.37%
III-C Cross-Modal Interaction
Our framework implements cross-modal fusion through a heterogeneous interaction module that projects textual and visual representations into a shared latent space. As shown in Figure 3, the proposed module is implemented via an alternating architecture composed of cross-modal interaction and relational graph learning. To enhance the representation of object geometric positions, we use an MLP (Multilayer Perceptron) to encode box vectors that are rotated to multiple viewpoints and then fused with the original object features as input. The interaction layers establish vision-language alignment through multi-head cross-attention mechanisms and the graph-relational layers encode structural relationships via graph attention networks. Noteworthy is the iterative application of the graph attention module, executed twice, with each iteration commenced by integrating geometric relation features into the input nodes. The vision-language cross-attention operation is formulated as:
This structure ensures that the graph nodes consistently contain linguistically informed cross-modal fusion content whenever the graph network engages in relational aggregation. The proposed relationship learning mechanism is constructed based on a multi-feature alignment framework that integrates semantic, spatial, and textual dimensions, which enhances the capability of the model to prioritize task-relevant nodes during iterative optimization processes selectively.
III-D Training
The training of our framework involves the following three parts: (i) a dual-branch 3D point cloud encoder, (iii) a text encoder with pre-trained BERT parameters initialized, and (ii) a cross-modal interaction module that consists of graph networks and cross-attention layers. To improve the robustness of the model, we employ a stochastic view selection strategy to the attribute-guided branch in the point cloud encoder. In the training stage, inputs are sampled from different viewpoints dynamically, while one representative view which has the most projected points is set as the input during inference. The composite loss function is formulated as a weighted combination:
where $\mathcal{L}_{ot}$ quantifies 3D object-text alignment using the metric defined in Section III-A1, $\mathcal{L}_{ref}$ addresses localization accuracy through predicted target probability distributions, $\mathcal{L}_{t}$ calculates a multi-class cross-entropy loss for text-based target classification, and $\mathcal{L}_{of}$ computes object-type classification loss based on fused features from the dual-branch point cloud encoder. The weighting coefficients $\lambda_{1},\lambda_{2},\lambda_{3}\in(0,1)$ balance task-specific objectives.
Figure 4: Visualization of our method on Nr3D and ScanRefer. The number under the text entry in the “Description” column indicates the number of target-class objects in the scene. The correctly predicted targets are surrounded by yellow bounding boxes, and the error predicted results are marked with red bounding boxes.
Figure 5: Visualization of the results of ablation experiments. The “baseline” denotes the complete AS3D method we propose. The last three columns present the localization results of the model without the three important components. The incorrectly predicted 3D bounding boxes are highlighted in red.
IV Experiment
IV-A Datasets and Evaluation Metrics
IV-A1 Nr3D/Sr3D
IV-A2 ScanRefer
IV-A3 Evaluation Metrics
@@
IV-B Implementation Details
We use the identical 3-layer PointNet++ architecture employed by prior two-stage methods to sample 1024 points per object for encoding, yielding 768-dimensional object features. We map single-view attribute features extracted by CLIP-VIT-B/16 to a 256-dimensional space and couple them with geometric features in the third layer of the point cloud network. All multi-head attention layers in our model are configured with 8 heads. The loss weight coefficients $\lambda_{1}$ and $\lambda_{3}$ are set to 0.5, while $\lambda_{2}$ is set to 0.1. Our training protocol leverages a NVIDIA 3090 GPU for all models, employing the Adam optimizer [54] with a batch size of 16. Models are trained for 100 epochs on both the Nr3D and ScanRefer datasets, and for an extended 120 epochs on Sr3D. We initiate the learning rate as $5\times 10^{-4}$, employing a scheduled multiplicative decay factor of 0.65 at intervals of 10 epochs between the 30th and 80th training epochs.
IV-C Comparison with Others
IV-C1 Nr3D/Sr3D
As shown in Table I, our approach exceeds the state-of-the-art method on both the Nr3D and Sr3D datasets by 2.0% and 2.1%, respectively. The improvements of 3.8% and 2.4% on hard samples indicate that our model can better learn object relationships in scenes with more distractors. Regarding view-dependent samples of Sr3D, although our model is slightly inferior to MiKASA, its view fusion requires additional feature extraction from multiple rotated point clouds through data augmentation. In contrast, our framework only needs to extract features once from the point cloud in the original view, resulting in a significant reduction in computational cost.
IV-C2 ScanRefer
Table II shows the localization results (IoU0.5absent0.5$\geq$ 0.5≥ 0.5) of advanced methods on Scanrefer, encompassing both methods utilizing 2D information and those relying solely on point clouds. The column “2D Views” shows the number of 2D views per object (“Num”) and the pre-trained encoder networks for 2D image processing (“Enc”). When “Num” is “all”, it means that all 2D information in the scene views will be aggregated.
@@ IV-C2 ScanRefer
Current methods mainly employ the pre-trained ENet segmentation model to extract pixel-level features from every 2D view, which are then projected onto 3D points. While this approach enhances 2D features, it comes with significant computational costs and redundant visual data due to the high overlap among adjacent views. In contrast, our method leverages CLIP for object-level feature fusion, requiring only one image for feature extraction per object. Compared to M3DRef-CLIP, which also uses CLIP to encode 2D objects, our AS3D surpasses it with fewer images. The 4.0% improvement on “Multiple” demonstrates that our approach is more effective at understanding positional relationships within the text, thereby enabling the distinction between similar objects. However, it’s important to note that our localization accuracy decreases in situations without distractors (“Unique”). This is due to our dependence on a training-free, class-agnostic segmentation model during the initial phase. In contrast, established methods use class-discriminative segmentation models, like PointGroup and VoteNet, which are trained with object classification. For targets that lack co-occurring objects in the same category, the semantic category priors alone are enough to localize the single entity.
IV-D Visualization Analysis
@@
Figure 4 presents visualization results of our method on the Nr3D and ScanRefer datasets. Both datasets are manually annotated with free-form textual descriptions, where sentences in ScanRefer are typically longer and contain more attribute information. The localization of Nr3D is directly based on GT boxes, while ScanRefer is based on predicted boxes from the segmentation model. The results demonstrate that our proposed method achieves precise localization accuracy in challenging scenes containing over two distractors. Nevertheless, upon analyzing the failure cases, we identify two primary shortcomings of our approach: decreased sensitivity to variations in distance and a drop in performance when dealing with perspective-dependent descriptions.
TABLE III: Ablation studies of key components on Nr3D and ScanRefer
@@
Overall Hard Easy V-dep V-indep Overall Unique Multiple
baseline 67.2% 61.4% 73.3% 65.7% 68.0% 45.37% 64.08% 40.87%
w/o 2D-assisted alignment 65.3% 59.5% 71.7% 64.9% 65.6% 44.86% 63.74% 40.32%
w/o attribute-guided branch 58.7% 53.4% 64.2% 57.8% 59.1% 42.50% 61.28% 37.97%
w/o graph attention 65.0% 58.6% 71.7% 64.3% 65.4% 44.73% 63.95% 40.10%
Figure 6: Visualization of graph attention results. The third column shows the heat maps of the attention weights of the last graph attention layer, where only the top 10 responsive node pairs are retained to represent object scores. In the second column, Orange ellipses demarcate possible references related to a descriptive relationship. Green bounding boxes denote GT targets, while the predictions of our model are color-coded with yellow for correct and red for incorrect.
IV-E Ablation Studies
IV-E1 Analysis of Key Components
Table III presents ablation studies for the Nr3D and ScanRefer datasets, showcasing the effectiveness of three key components: the 2D-assisted vision-language alignment (Section III-A1), the attribute-guided branch of the point cloud encoder (Section III-A2), and the graph attention module based on the semantic-spatial scene graph (Section III-B). The standardized bounding box annotations of Nr3D aim to separate detector performance from the evaluation metric, thereby enabling a direct assessment of referential language comprehension capabilities. The metrics of ScanRefer involve the accuracy of the 3D bounding box, therefore, in addition to multi-modal comprehension, there is also a segmentation effect. “w/o graph attention” means removing all of the graph attention layers from the cross-modal interaction module. By comparing the results of “w/o 2D-assisted alignment” to the baseline AS3D, it is evident that 3D features trained using a 2D pre-trained multi-modal model align more easily with text. The slight improvement in viewpoint dependency challenges arises because 2D vision does not assist in perceiving viewpoint information. The results of “w/o graph attention” show that the relation learning of the semantic-spatial scene graph has significantly improved 3D localization based on referential descriptions, especially on hard samples with complex environments. The object attribute-guided feature encoding obtained from CLIP demonstrates the most notable performance improvement across all evaluation metrics. This underscores the effectiveness of our proposed dual-branch point cloud encoding network in significantly improving the 3D visual understanding capabilities of the multi-modal grounding model.
The visualization results in Figure 5 demonstrate the role of each component in 3D visual grounding for multi-modal understanding. The 2D-assisted alignment module and the attribute-guided branch setting focus more on the inherent object features of attributes and class-aware semantics. The absence of these two components may lead to erroneous perception, such as misidentification of non-target object categories in the second-row example or ambiguity in encoding additional attributes of the referenced object (“chair”) in the last-row case. Relation learning based on scene graphs primarily manifests in the comprehension of referential relationships within textual descriptions. Particularly when the target is close to homogeneous distractors (as exemplified in the fourth-row case), the graph attention mechanism demonstrates enhanced capability in distinguishing the target object through leveraging complex spatial relationships. Figure 6 illustrates the maximum attention scores assigned to nodes during feature aggregation in the final graph attention layer. It can be seen that for accurately predicted samples, the relation graph assigns higher attention scores to target or referred objects that match the textual description. Notably, in the last row, although the target object receives a relatively higher attention score, the aggregated features do not offer a competitive advantage in the subsequent text alignment processing, ultimately resulting in incorrect localization.
TABLE IV: Ablation studies of 3D object-text Loss on Nr3D
  $PI^{*}$ $TT^{*}$ $IC$ Overall Hard Easy V-dep V-indep
@@ IV-E1 Analysis of Key Components
TABLE V: Ablation studies of object encoder branch structure on Nr3D

### v3

@@ 3.2 Language-Guided Scene Graph
In the 3D visual grounding task, accurately distinguishing between identical objects mainly depends on their positional context and referential relationships, as described in the given description. To tackle this issue, we introduce relational graph learning. The scene graph focuses only on objects semantically related to the referential description. This reduces noise and improves reasoning efficiency. To enhance spatial reasoning capacities, we integrate object-level spatial priors with global spatial encoding [11], incorporating them into the object features prior to graph attention operations. We also employ the rotated bounding box encoding paradigm [7] to facilitate multi-view spatial comprehension via decoupled shared point cloud features.
3.2.1 Graph Generation via Semantic Consistency
@@
In complex scenes, the target description often entails multiple referential relationships. The innate capability of scene graphs to capture dependencies among various nodes aids in effectively processing intricate scene compositions. Prior studies [14, 16, 45] either construct fully-connected scene graphs or establish connections with neighboring nodes, implicitly learning the node and text embedding relations. However, many positional descriptions are not dependent on proximity, and a large number of objects can introduce uncertainties and extra cost in node aggregation. We propose an explicit construction of graph node connections using 3D-language matching about object semantics. First, we match 3D feature representations to class-aware prompts as described in Section 3.1.1. Then, we utilize the SpaCy lemmatization module to standardize the description text by converting words to their base forms, thereby significantly improving terminological consistency across different morphological variants. Finally, we use lexemes from the same semantic category in the modified description text to identify related objects. These matched objects, considered potential targets or referred objects in localization cues, are then connected within the graph. It ensures the graph is not a generic scene representation, but a task-specific reasoning graph conditioned on the input language.
3.2.2 Relational Learning via Graph Attention
@@
The graph nodes are input into a graph attention layer to facilitate multi-relational interactions, where feature representations are enhanced by aggregating heterogeneous relation information from neighboring nodes. The $i$-th node value is updated through a multi-head attention mechanism:
@@
where $\mathcal{N}(i)$ denotes the set of neighboring nodes of vertex $i$, $A_{ij}\in\{0,1\}$ denotes the original adjacency relation between nodes $i$ and $j$, $\alpha_{ij}^{k}\in(0,1]$ captures the normalized attention weight from node $j$ to $i$ through the $k$-th attention head. $W^{k}$ is the learnable projection matrix for the $k$-th head, $\sigma(\cdot)$ specifies the non-linear activation function.
3.3 Cross-Modal Interaction
Our framework implements cross-modal fusion through a heterogeneous interaction module that projects textual and visual representations into a shared latent space. As shown in Figure Fig. 3, the proposed module is implemented via an alternating architecture composed of cross-modal interaction and relational graph learning. To enhance the representation of object geometric positions, we use an MLP (Multilayer Perceptron) to encode box vectors that are rotated to multiple viewpoints and then fused with the original object features as input. The interaction layers establish vision-language alignment through multi-head cross-attention mechanisms and the graph-relational layers encode structural relations via graph attention networks. Noteworthy is the iterative application of the graph attention module, executed twice, with each iteration commenced by integrating geometric relational features into the input nodes. The vision-language cross-attention operation is formulated as:
This structure ensures that the graph nodes consistently contain linguistically informed cross-modal fusion content whenever the graph network engages in relational aggregation. The proposed relational learning mechanism is constructed based on a multi-feature alignment framework that integrates semantic, spatial, and textual dimensions, which enhances the capability of the model to prioritize task-relevant nodes during iterative optimization processes selectively.
3.4 Training
The training of our framework involves the following three parts: (i) a dual-branch 3D point cloud encoder, (ii) a text encoder with pre-trained BERT parameters initialized, and (iii) a cross-modal interaction module that consists of graph networks and cross-attention layers. To improve the robustness of the model, we employ a stochastic view selection strategy to the 2D-enhanced branch in the object encoder. In the training stage, inputs are sampled from different viewpoints dynamically, while one representative view which has the most projected points is set as the input during inference. The composite loss function is formulated as a weighted combination:
where $\mathcal{L}_{ot}$ quantifies 3D object-text alignment using the metric defined in Section 3.1.1, $\mathcal{L}_{ref}$ addresses localization accuracy through predicted target probability distributions, $\mathcal{L}_{t}$ calculates a multi-class cross-entropy loss for text-based target classification, and $\mathcal{L}_{of}$ computes object-type classification loss based on fused features from the dual-branch point cloud encoder. The weighting coefficients $\lambda_{1},\lambda_{2},\lambda_{3}\in(0,1)$ balance task-specific objectives.
4 Experiment
4.1 Datasets and Evaluation Metrics
4.1.1 Nr3D/Sr3D
4.1.2 ScanRefer
4.1.3 Evaluation Metrics
[…]
4.2 Implementation Details
We use the identical 3-layer PointNet++ architecture employed by prior two-stage methods to sample 1024 points per object for encoding, yielding 768-dimensional object features. We map single-view features extracted by CLIP-VIT-B/16 to a 256-dimensional space and couple them with geometric features in the third layer of the point cloud network. All multi-head attention layers in our model are configured with 8 heads. The loss weight coefficients $\lambda_{1}$ and $\lambda_{3}$ are set to 0.5, while $\lambda_{2}$ is set to 0.1. Our training protocol leverages a NVIDIA 3090 GPU for all models, employing the Adam optimizer with a batch size of 16. Models are trained for 100 epochs on both the Nr3D and ScanRefer datasets, and for an extended 120 epochs on Sr3D. We initiate the learning rate as $5\times 10^{-4}$, employing a scheduled multiplicative decay factor of 0.65 at intervals of 10 epochs between the 30th and 80th training epochs.
4.3 Comparison with Others
4.3.1 Nr3D/Sr3D
Table 1: Comparison with other recent methods on Nr3D and Sr3D
@@ 4.3.1 Nr3D/Sr3D
LSVG (ours) - 67.2% 61.4% 73.3% 65.7% 68.0% 76.0% 71.7% 77.8% 67.2% 76.4%
vs. MiKASA [11] $\uparrow$2.8% $\uparrow$2.0% $\uparrow$3.6% $\uparrow$0.3% $\uparrow$4.0% $\uparrow$0.8% $\uparrow$4.4% $\downarrow$0.8% $\downarrow$3.2% $\uparrow$1.0%
vs. MA2TransVG [10] $\uparrow$2.0% $\uparrow$3.8% $\uparrow$2.1% $\uparrow$3.2% $\uparrow$2.6% $\uparrow$2.1% $\uparrow$2.4% $\uparrow$1.8% $\uparrow$2.7% $\uparrow$2.6%
As shown in Table 1, our approach exceeds the state-of-the-art method on both the Nr3D and Sr3D datasets by 2.0% and 2.1%, respectively. The improvements of 3.8% and 2.4% on hard samples indicate that our model can better learn object relationships in scenes with more distractors. Regarding view-dependent samples of Sr3D, although our model is slightly inferior to MiKASA, its view fusion requires additional feature extraction from multiple rotated point clouds through data augmentation. In contrast, our framework only needs to extract features once from the point cloud in the original view, resulting in a significant reduction in computational cost.
4.3.2 ScanRefer
Table 2: Comparison with other recent methods on Scanrefer.
  Methods 2D Views Multiple Overall
View Num Encoder
TGNN [16] - - 23.18% 29.70%
InstanceRefer [17] - - 24.77% 32.93%
MVT [7] - - 25.26% 33.26%
ViewRefer [8] - - 26.50% 33.66%
BUTD-DETR [48] - - 32.81% 37.05%
EDA [47] - - 37.64% 42.26%
MCLN [49] - - 38.41% 42.64%
ScanRefer [19] all ENet 21.11% 27.40%
3DVG-Trans [20] all ENet 28.42% 34.67%
3D-SPS [21] all ENet 29.82% 36.98%
3DJCG [22] all ENet 30.82% 37.33%
D3Net [23] all ENet 30.05% 37.87%
3D-LP [50] all ENet 33.41% 39.46%
3DVLP [44] all ENet 33.40% 40.51%
MA2TransVG [10] all ENet 41.4% 45.7%
M3DRef-CLIP [24] 3 CLIP 36.8% 44.7%
LSVG (ours) 1 CLIP 40.87% 45.37%
@@ 4.3.2 ScanRefer
Current methods mainly employ the pre-trained ENet segmentation model to extract pixel-level features from every 2D view, which are then projected onto 3D points. While this approach enhances 2D features, it comes with significant computational costs and redundant visual data due to the high overlap among adjacent views. In contrast, our method leverages CLIP for object-level feature fusion, requiring only one image for feature extraction per object. Compared to M3DRef-CLIP, which also uses CLIP to encode 2D objects, our LSVG surpasses it with fewer images. The 4.0% improvement on “Multiple” demonstrates that our approach is more effective at understanding positional relationships within the text, thereby enabling the distinction between similar objects.
4.4 Visualization Analysis
@@
Figure Fig. 4 presents visualization results of our method on the Nr3D and ScanRefer datasets. Both datasets are manually annotated with free-form textual descriptions, where sentences in ScanRefer are typically longer and contain more attribute information. The localization of Nr3D is directly based on GT boxes, while ScanRefer is based on predicted boxes from the segmentation model. The results demonstrate that our proposed method achieves precise localization accuracy in challenging scenes containing over two distractors. Nevertheless, upon analyzing the failure cases, we identify two primary shortcomings of our approach: decreased sensitivity to variations in distance and a drop in performance when dealing with perspective-dependent descriptions.
Fig. 4: Visualization of our method on Nr3D and ScanRefer. The number under the text entry in the “Description” column indicates the number of target-class objects in the scene. Yellow bounding boxes surround the correctly predicted targets, and the error predicted results are marked with red bounding boxes.
Table 3: Ablation studies of key components on Nr3D and ScanRefer
@@
Overall Hard V-dep Overall Multiple
baseline 67.2% 61.4% 65.7% 45.37% 40.87%
w/o 2D supervision 65.3% 59.5% 64.9% 44.86% 40.32%
w/o 2D enhancement 58.7% 53.4% 57.8% 42.50% 37.97%
w/o graph learning 65.0% 58.6% 64.3% 44.73% 40.10%
4.5 Ablation Studies
4.5.1 Analysis of Key Components
Table 4: Ablation studies of 3D object-text loss on Nr3D
  $PI^{*}$ $TT^{*}$ ICIC Overall Hard Easy V-dep V-indep
@@ 4.5.1 Analysis of Key Components
Table 3 presents ablation studies for the Nr3D and ScanRefer datasets, showcasing the effectiveness of three key components: the 2D-supervised 3D-language alignment (Section 3.1.1), the 2D-enhanced encoding branch (Section 3.1.2), and the graph learning module based on the language-guided scene graph (Section 3.2). The standardized bounding box annotations of Nr3D aim to separate detector performance from the evaluation metric, thereby enabling a direct assessment of referential language comprehension capabilities. The metrics of ScanRefer involve the accuracy of the 3D bounding box. Therefore, in addition to multi-modal comprehension, there is also a segmentation effect. “w/o graph learning” means removing all of the graph attention layers from the cross-modal interaction module. By comparing the results of “w/o 2D supervision” to the baseline LSVG, it is evident that 3D features trained using a 2D pre-trained multi-modal model align more easily with text. The slight improvement in viewpoint dependency challenges arises because 2D vision does not assist in perceiving viewpoint information. The results of “w/o graph learning” show that relational graph learning has significantly improved 3D localization based on referential descriptions, especially on hard samples with complex environments. The 2D-enhanced encoding obtained from CLIP demonstrates the most notable performance improvement across all evaluation metrics (“w/o 2D enhancement”). This underscores the effectiveness of our proposed dual-branch point cloud encoding network in significantly improving the 3D visual understanding capabilities of the multi-modal grounding model.
Fig. 5: Visualization of the results of ablation experiments. The “baseline” denotes the complete LSVG method we propose. The last three columns present the localization results of the model without the three important components. The incorrectly predicted 3D bounding boxes are highlighted in red.
Fig. 6: Visualization of graph attention results. The third column shows the heat maps of the attention weights of the last graph attention layer, where only the top 10 responsive node pairs are retained to represent object scores. In the second column, Orange ellipses demarcate possible references related to a descriptive relationship. Green bounding boxes denote GT targets, while the predictions of our model are color-coded with yellow for correct and red for incorrect.
The visualization results in Figure Fig. 5 demonstrate the role of each component in 3D visual grounding for multi-modal understanding. The 2D-assisted modules focus more on the inherent object attributes and class-aware semantics. The absence of these two components may lead to erroneous perception, such as misidentification of non-target object categories in the second-row example or ambiguity in encoding additional attributes of the referred object (“chair”) in the last-row case. Relational learning based on scene graphs primarily manifests in the comprehension of referential relationships within textual descriptions. Particularly when the target is close to homogeneous distractors (as exemplified in the fourth-row case), the graph attention mechanism demonstrates enhanced capability in distinguishing the target object through leveraging complex spatial relationships. Figure Fig. 6 illustrates the maximum attention scores assigned to nodes during feature aggregation in the final graph attention layer. It can be seen that for accurately predicted samples, the relational graph assigns higher attention scores to target or referred objects that match the textual description. Notably, in the last row, although the target object receives a relatively higher attention score, the aggregated features do not offer a competitive advantage in the subsequent text alignment processing, ultimately resulting in incorrect localization.
4.5.2 Losses of 2D-Assisted Training
In Section 3.1.1, we employ various contrastive losses to define the 3D object-text alignment loss. The losses that leverage 2D pre-trained multi-modal features play an important role in 3D vision-language learning. Table 4 presents the quantitative results of the impact of various loss functions on localization performance. $PI^{*}$” denotes the contrastive loss between the point cloud features and CLIP-extracted 2D features. $TT^{*}$” represents the contrastive loss between text features extracted by our framework and CLIP-generated textual embeddings. “ICIC” refers to the interactive contrastive loss that aligns 3D and 2D visual features with the interchangeable corresponding text features. The tabular data demonstrates that both text contrastive and interaction contrastive losses substantially improve 3D multi-modal learning. This suggests that incorporating 2D-assisted training can help mitigate the issue of cross-modality understanding in situations where 3D data is limited.
4.5.3 Structure of Object Encoder
Table 5: Ablation studies of object encoder branch structure on Nr3D